In a distribution for which the mean is 25 and the standard deviation is 5, what percentage of all scores occur at 30 or above?
Answer by fractalier(6550)

Scores 30 and above occur at greater than or equal to one standard deviation above the mean.
We know 68.2% of a sample fits within one standard deviation. Thus the part more than 1 s.d. is (1/2)(100% - 68.2%) = (1/2)(31.8%) = 15.9%